A 228/-30 volt constant potential generator delivers load to an impedance. Assuming the reactance only to be variable and the line impedance is (6 + j4.5) ohms, calculate for maximum power transfer: a) the load impedance, b) the current, and c) the load power and power factor. NEEDS SOLUTION


1
Expert's answer
2021-04-22T07:22:56-0400

a) For maximum power transfer the load impedance "=6-j4.5"


b) The current "=228\/12= 19A"


c) The load power"=I^2R=19^2\times6=2166W"


The load power factor is "Cos(tan^{-1}(4.5/6) )=cos36.86\degree=0.8lead"